f(6) = 3,f(9) = 15
Answer by stanbon(75887)

f(6) = 3,f(9) = 15
---
slope = (15-3)(9-6) = 12/3 = 4
----
Equation Form:
3 = 4*6 + b
b = -21
-------
Equation:
f(x) = 4x-21
==================
Cheers,
Stan H.
==================